OpenStudy (anonymous):

show all work. Find the product: (1/7 x^2 - 2/5) (1/2x - 1/5)

OpenStudy (anonymous):

\[\LARGE \left(\frac17 \cdot x^2-\frac25\right)\left(\frac12 \cdot x-\frac 15\right)\] this ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well I think it is. Let me go grab the screen image.

OpenStudy (anonymous):

\[\LARGE \left(\frac{x^2}{7} -\frac25\right)\left(\frac{x}{2}-\frac 15\right)\] do you know to find common denominator ?

OpenStudy (anonymous):

No. But this is what the problem looks like

OpenStudy (anonymous):

this is it ... I'll ask you again, do you know to find common denominator ? :)

OpenStudy (anonymous):

No. I am terrible at math.

OpenStudy (anonymous):

hmm.... I can do it for you, but I can't keep teaching you each step, it's going to take long time ! ... would you like me to do it or not ?

OpenStudy (anonymous):

yes please. I normally have my friend to walk me through step by step but she is not on right now. :/

OpenStudy (anonymous):

\[\LARGE \left(\frac{x^2}{7}-\frac25\right)\left(\frac x2- \frac 15\right)=\left(\frac{5x^2}{35}-\frac{14}{35}\right)\left(\frac{5x}{10}- \frac {2}{10}\right)\] \[\LARGE =\left(\frac{5x^2-14}{35}\right)\left(\frac{5x-2}{10} \right)=\] \[\LARGE =\left(\frac{25x^3 -10x^2 -60x+28}{350 }\right) \]

OpenStudy (anonymous):

thank you so much!